OpenStudy (anonymous):

what are the integral of dx/(e^x +1) and dx/((e^x + 1)e^2x)

OpenStudy (jamesj):

Notice that \[ \frac{1}{e^x+1} = \frac{e^x + 1}{e^x+1} - \frac{e^x}{e^x+1} \]

OpenStudy (anonymous):

but im no used to these seperation of fractions. really how you know when to separate fractions?

OpenStudy (jamesj):

I'm not using partial fractions; I'm just trying to write 1/(e^x+1) as the sum of functions we can integrate. Clearly the first term (e^x+1)/(e^x+1) = 1, which you can integrate. Look now at the second term and see why you can integrate that.

OpenStudy (anonymous):

i know the answer now thnx. but im not sure if i can do it in tst ^.^

OpenStudy (jamesj):

That's why you practice and see different examples, to help build up your personal tool kit.